# VidCRAFT3: Camera, Object, and Lighting Control for Image-to-Video Generation

### v3

@@ 3.2 Model Architecture
Camera Motion Control via Point Cloud Rendering. To achieve precise camera motion control in I2V, VidCRAFT3 leverages the Image2Cloud module, which reconstructs a high-quality 3D point cloud of the scene from a single reference image. Specifically, we employ DUSt3R, an unconstrained stereo 3D reconstruction model, to generate a 3D point cloud. Given a reference image $I_{\text{ref}}$, DUSt3R performs monocular or binocular reconstruction via point regression, followed by global alignment to ensure multi-view consistency: $\mathcal{P}=\text{DUSt3R}(I_{\text{ref}})$. The reconstructed point cloud provides explicit 3D geometry, enabling accurate rendering of the scene from arbitrary camera motion trajectory $E=\{E_{1},E_{2},\ldots,E_{F}\}$, where $F$ is the length of trajectory. Given a user-defined trajectory $E$, the point cloud rendering at frame $t$ is computed as $R_{t}=\pi(\mathcal{P},E_{t})$, where $\pi(\cdot)$ is the differentiable rendering function. However, due to the limitations of point cloud representation and the sparse 3D cues from a single image, the point cloud renderings may exhibit artifacts such as missing regions, occlusions, and geometric distortions. To address this, VidCRAFT3 integrates point cloud renderings as a input of VDM, which refines the coarse renderings to generate high-quality and temporally consistent video frames. This combination of explicit 3D geometry and generative refinement ensures both accurate camera control and realistic video synthesis.
@@
Object Motion Control through Trajectory Learning. Object motion in VidCRAFT3 is controlled through sparse spatial trajectories. For up to $N$ objects in a $F$-frames video, each trajectory is defined as a sequence of 2D pixel coordinates $\mathcal{T}=\{\mathbf{s}_{n}^{f}=(x_{n}^{f},y_{n}^{f})\mid n\in[1,N],\,f\in[1,F]\}$, where $\mathbf{s}_{n}^{f}$ denotes the position of the $n$-th object in frame $f$. To model motion dynamics, we compute inter-frame optical flow vectors between consecutive frames. For each trajectory point $\mathbf{s}_{n}^{f}$, the displacement vector $\mathbf{v}_{n}^{f}$ is calculated as $\mathbf{v}_{n}^{f}=\mathbf{s}_{n}^{f+1}-\mathbf{s}_{n}^{f}=(x_{n}^{f+1}-x_{n}^{f},\,y_{n}^{f+1}-y_{n}^{f}).$ These sparse motion vectors are then projected into a per-frame optical flow map $\mathcal{V}^{f}\in\mathbb{R}^{H\times W\times 2}$, where $H$ and $W$ denote spatial dimensions. The mapping is formalized as
@@
with the first frame’s flow initialized as $\mathcal{V}^{1}(x,y)=(0,0),\forall(x,y)$. The full spatiotemporal flow tensor $\mathcal{V}\in\mathbb{R}^{F\times H\times W\times 2}$ is subsequently processed through Gaussian smoothing to obtain a dense motion representation $\tilde{\mathcal{V}}$. The ObjMotionNet is a neural network composed of multiple convolutional layers and downsampling operations, designed to extract multi-scale motion features from $\tilde{\mathcal{V}}$. Inspired by T2I-Adapter [33], ObjMotionNet injects multi-scale motion features exclusively into the UNet Encoder. This balances precise motion control with video quality, as the encoder handles structure while the decoder refines details, ensuring accurate guidance without compromising output quality.
Figure 3: Construction pipeline of the Object Motion Control Dataset.
Lighting Direction Control with Spatial Triple-Attention Transformer. The lighting direction is represented as a per-frame 3D vector $L=(l_{x},l_{y},l_{z})$, describing the orientation of light source in Cartesian coordinates. To effectively encode this directional information into a high-dimensional feature space, we employ Spherical Harmonic (SH) Encoding. SH encoding captures the angular characteristics of the lighting using basis functions up to degree 4, resulting in 16 coefficients. The resulting SH-encoded vector $L_{\text{SH}}\in\mathbb{R}^{16}$ is projected into the feature space of the UNet using a multi-layer perceptron (MLP). $\mathbf{E}_{\text{light}}=\text{MLP}(L_{\text{SH}})$, where $\mathbf{E}_{\text{light}}$ is a lighting embedding aligned with the dimensionality of text embedding.
@@
where $Q$ (Query) comes from the self-attention output of the UNet, $K$ (Key) and $V$ (Value) are derived from $\mathbf{E}_{\text{light}}$. The outputs of three cross-attention are summed to produce a fused feature representation $\mathbf{O}=\mathbf{O}_{\text{image}}+\mathbf{O}_{\text{text}}+\mathbf{O}_{\text{light}}$ where $\mathbf{O}_{\text{image}}$, $\mathbf{O}_{\text{text}}$, and $\mathbf{O}_{\text{light}}$ are the outputs of the image, text, and lighting cross-attention modules, respectively. This novel mechanism ensures the generated videos maintain consistency across lighting, text, and image conditions.
Figure 4: Construction pipeline of the VLD Dataset
3.3 Dataset Construction
@@
Due to the lack of datasets with annotations for camera motion trajectory, object motion trajectories, and lighting directions, we construct three specialized datasets. All datasets consist of 25-frame video clips with a spatial resolution of $320\times 512$ pixels.
@@
Camera Motion Control Dataset. We constructe this dataset from RealEstate10K [75], creating 62,000 clips with smooth camera transitions. Using DUSt3R, we generated globally aligned point clouds and rendered the first frame’s point cloud along the ground-truth camera trajectory. Since RealEstate10K lacks captions, we uniformly sampled 4 frames per clip and generated descriptions using Qwen2-VL-7B-Instruct [53]. Despite its focus on indoor scenes, the dataset provides diverse camera trajectories for fine-grained motion control.
@@
Object Motion Control Dataset. We construct this dataset from WebVid-10M [2], consisting of 60,000 video clips. The dataset construction pipeline, illustrated in Fig. 3, consists of the following five steps: (1) Clip Filtering: Clips with abrupt scene changes are removed using PySceneDetect, and sequences with temporal intervals (1–16 frames) are sampled. To retain only clips with significant motion, optical flow is computed via MemFlow [10], and the bottom 25% of clips with low motion scores are filtered out. (2) Dense Trajectory Generation: For each filtered clip, CoTrackerV3 [24] generates dense trajectories on a $16\times 16$ grid, capturing pixel-level motion dynamics across frames. (3) Trajectory Filtering by Length: Trajectories shorter than the average length are discarded to ensure meaningful object motion. (4) Sparse Trajectory Sampling: From the filtered dense trajectories, 1–8 sparse trajectories are sampled with probability proportional to their length. (5) Optical Flow Smoothing: Optical flow between adjacent frames is computed to encode motion direction and intensity. Finally, a Gaussian filter is applied to smooth the sparse trajectory matrix, ensuring stable training.
@@
Lighting Direction Control Dataset. Since it is challenging to control a camera to follow the same trajectory under varying lighting conditions in real-world scenarios, collecting such datasets is both extremely difficult and costly. To address this, we introduce the VideoLightingDirection (VLD) Dataset, a novel synthetic dataset designed for modeling complex light-object interactions. The VLD Dataset comprises 57,600 synthetic videos rendered using Blender, featuring 3,600 realistic scenes, each with 16 videos captured under distinct lighting conditions. As illustrated in Fig. 4, the dataset construction pipeline consists of the following four steps: (1) Scene Creation: We designed two scene types: Haven scenes and BOP scenes. Haven scenes places 3D models from Poly Haven at the center of HDR environments, while BOP scenes randomly places BOP models in a six-plane textured room to simulate indirect lighting. Models and environments can be randomly combined for diversity. See supplementary materials for BOP scenes examples. (2) Camera Trajectory Sampling: We sample the starting point of the camera trajectory on a spherical region (radius: 0.7–1.3 meters) centered around the 3D models. A smooth trajectory is then randomly generated around the model, ensuring that the camera always facing the center of the model. (3) Lighting Direction Sampling: To enhance the lighting effect, we reduce the HDR environment intensity by 40%. We uniformly sample 16 points on a hemisphere centered on the model, with the base surface normal aligned with the camera’s viewing direction. Each sampled point serves as the position of a 2kW spotlight (radius: 1), oriented toward the model’s center. The light source position vectors are mapped to all camera coordinate systems and normalized to compute the lighting direction for each view. (4) Rendering and Annotation: Using Blender Cycles, we render each scene under 16 different lighting conditions while maintaining a consistent camera trajectory. Each frame is annotated with the corresponding camera pose and lighting direction, providing precise control signals for training.
3.4 Training Strategy
To enable simultaneous control over camera motion, object motion, and lighting direction, we adopt a three-stage progressive training strategy, designed to address the absence of datasets annotated with all three control signals.
@@ 3.4 Training Strategy
Stage 1: Camera Motion Control Training. The training begins by initializing the model with DynamiCrafter pre-trained weights. The model is then fine-tuned on camera motion control dataset for 40,000 iterations, optimizing the entire UNet to align VDM with camera motion. This stage establishes robust 3D scene understanding by integrating point cloud renderings, ensuring precise global camera movements while maintaining temporal consistency.
@@
Stage 2: Dense Object Trajectories and Lighting Mixed Fine-tuning. We combine the Object Motion Control Dataset and VLD Datasets to create a comprehensive dataset annotated with camera motion, object motion, and lighting direction. This hybrid dataset enhances the ability of model to learn joint control over these three conditions. Dense object trajectories are incorporated to provide rich motion details, accelerating model convergence. To retain the camera control capabilities from stage 1, the temporal layers of the UNet remain frozen, while the spatial layers and newly introduced components, including ObjMotionNet, lighting cross-attention, and the MLP for lighting direction projection, are optimized for 20,000 iterations. This stage enables the model to simultaneously control all three conditions while ensuring global camera alignment.
@@
Stage 3: Sparse Object Trajectories and Lighting Mixed Fine-tuning. The same hybrid dataset from Stage 2 is reused, but dense trajectories are replaced with sparse trajectories to simulate real-world user interactions. Fine-tuning follows the same parameters as Stage 2, with 20,000 iterations. This stage forces the model to infer complex motion patterns from limited trajectory data, and the progressive shift from dense to sparse supervision enhances its ability to generalize to practical scenarios. By progressively learning camera motion, object motion, and lighting direction controls across stages while strategically freezing layers to retain prior knowledge, the training strategy ensures that VidCRAFT3 achieves fine-grained, synergistic control over all three elements, even in the absence of fully annotated multi-task training data.
4 Experiments
Figure 5: Qualitative comparisons with SOTA methods on RealEstate10K.
Figure 6: Qualitative comparisons with SOTA methods on WebVid-10M.
4.1 Experimental Setup

### v4

@@ III-C Model Architecture
Camera Motion Control via Point Cloud Rendering. Directly injecting camera parameters into the UNet learns an implicit mapping of camera poses to dynamics [7, 6], often causing coarse control, weak 3D consistency, and poor generalization to unseen trajectories. Inspired by [41, 42, 44], VidCRAFT3 leverages the Image2Cloud module, which reconstructs a high-quality 3D point cloud of the scene from a single reference image to provide explicit 3D priors, and combines it with a VDM for photorealistic refinement. Specifically, we employ DUSt3R, an unconstrained stereo 3D reconstruction model, to generate a 3D point cloud. Given a reference image $I_{\text{ref}}$, DUSt3R performs monocular or binocular reconstruction via point regression, followed by global alignment to ensure multi-view consistency: $\mathcal{P}=\text{DUSt3R}(I_{\text{ref}})$. The reconstructed point cloud provides explicit 3D geometry, enabling accurate rendering of the scene from arbitrary camera trajectory. Given a user-defined camera trajectory $E=\{E_{f}\}_{f=1}^{F}$ with $E_{f}\in SE(3)$, the point cloud rendering at frame $t$ is computed as $R_{f}=\pi(\mathcal{P},E_{f})$, where $\pi(\cdot)$ is the differentiable rendering function. We replace the first point cloud rendering with the reference image, i.e., $R_{1}=I_{\text{ref}}$, which enforces an exact first-frame match and reduces drift from point-cloud noise and camera pose. Concretely, we use the per-frame rendering as the camera-conditioning input, i.e., $c_{\text{cam}}=\{R_{f}\}_{f=1}^{F}$. Following DynamiCrafter, we encode the point cloud rendering with the VAE, sample noise $\epsilon_{t}\!\sim\!\mathcal{N}(0,I)$, and concatenate it channel-wise to form the model input in latent space. Due to the limitations of point cloud representation and the sparse 3D cues from a single image, the point cloud renderings may exhibit artifacts such as missing regions, occlusions, and geometric distortions. To address this, VidCRAFT3 integrates point cloud renderings as an input to the VDM, which refines the coarse renderings to generate high-quality and temporally consistent video frames. This combination of explicit 3D geometry and VDM ensures both accurate camera control and realistic video synthesis.
@@
Object Motion Control through Trajectory Learning. Object motion in VidCRAFT3 is controlled through sparse object trajectories that are drawn by the user on the reference image. For up to $N$ objects in an $F$-frames video, each trajectory is defined as a sequence of 2D pixel coordinates $\mathcal{T}=\{\mathbf{s}_{n}^{f}=(x_{n}^{f},y_{n}^{f})\}$, with $n\in\{1,\ldots,N\}$ and $f\in\{1,\ldots,F\}$. $\mathbf{s}_{n}^{f}$ denotes the position of the $n$-th object in frame $f$. To model motion dynamics, we compute inter-frame optical flow vectors. For each trajectory point $\mathbf{s}_{n}^{f}$, the displacement vector $\mathbf{v}_{n}^{f}$ is calculated as $\mathbf{v}_{n}^{f}=\mathbf{s}_{n}^{f+1}-\mathbf{s}_{n}^{f}=(x_{n}^{f+1}-x_{n}^{f},\,y_{n}^{f+1}-y_{n}^{f}),f\in\{1,\ldots,F-1\}.$ These sparse motion vectors are then projected onto a per-frame optical flow map $\mathcal{V}^{f}\in\mathbb{R}^{H\times W\times 2}$. The mapping is formalized as
@@
with the first frame’s flow initialized as $\mathcal{V}^{1}(x,y)=(0,0),\forall(x,y)$. The full spatiotemporal flow tensor $\mathcal{V}\in\mathbb{R}^{F\times H\times W\times 2}$ is subsequently processed through Gaussian smoothing to obtain a dense motion representation $\tilde{\mathcal{V}}$. The ObjMotionNet is a neural network composed of multiple convolutional layers and downsampling operations, designed to extract multi-scale motion features $c_{\text{obj}}$. Inspired by T2I-Adapter [99], ObjMotionNet injects multi-scale motion features exclusively into the UNet encoder. This balances precise motion control with video quality, as the encoder handles structure while the decoder refines details, ensuring accurate guidance without compromising output quality.
Lighting Direction Control with Spatial Triple-Attention Transformer. In VidCRAFT3, the lighting direction is specified by the user as a unit vector $L_{\text{ref}}=(l_{x},l_{y},l_{z})$ defined in the reference view’s camera coordinate system. To derive lighting directions for other viewpoints along the camera trajectory, we transform $L$ using the extrinsic parameters. First, we promote $L$ to a homogeneous coordinate $\tilde{\mathbf{p}}_{\mathrm{ref}}=[L^{\top},1]^{\top}$ and use the inverse of the reference extrinsic $E_{\mathrm{ref}}^{-1}=E_{1}^{-1}$ to obtain its world-space position: $\tilde{\mathbf{p}}_{\mathrm{world}}=E_{\mathrm{ref}}^{-1}\,\tilde{\mathbf{p}}_{\mathrm{ref}}$. For each subsequent frame, world-to-camera extrinsic $E_{f}$ maps $\tilde{\mathbf{p}}_{\mathrm{world}}$ into that camera’s coordinate system, and we normalize the resulting three-vector to obtain a unit lighting direction:
$\tilde{\mathbf{p}}_{f}=E_{f}\,\tilde{\mathbf{p}}_{\mathrm{world}},\quad L_{f}=\frac{[\tilde{\mathbf{p}}_{f}]_{1:3}}{\|[\tilde{\mathbf{p}}_{f}]_{1:3}\|}.$
The sequence $L=\{L_{f}\}_{f=1}^{F}$ thus forms the per-frame lighting directions used in our model. To effectively encode this directional information into a high-dimensional feature space, we employ Spherical Harmonic (SH) Encoding. SH encoding captures the angular characteristics of the lighting using basis functions up to degree 4, resulting in 16 coefficients. The resulting SH-encoded vector $L_{\text{SH}}\in\mathbb{R}^{F\times 16}$ is projected into the feature space of the UNet using a multi-layer perceptron (MLP). $c_{\text{light}}=\text{MLP}(L_{\text{SH}})$, where $c_{\text{light}}$ is a lighting embedding aligned with the dimensionality of the text embedding.
@@
where $Q$ (Query) comes from the self-attention output of the UNet, $K$ (Key) and $V$ (Value) are derived from $c_{\text{light}}$. The outputs of the three cross-attention modules are summed to produce a fused feature representation $\mathbf{O}=\mathbf{O}_{\text{image}}+\mathbf{O}_{\text{text}}+\mathbf{O}_{\text{light}}$, where $\mathbf{O}_{\text{image}}$, $\mathbf{O}_{\text{text}}$, and $\mathbf{O}_{\text{light}}$ are the outputs of the image, text, and lighting cross-attention modules, respectively. This novel mechanism ensures the generated videos maintain consistency across lighting, text, and image conditions.
III-D Dataset Construction
@@
Due to the lack of datasets annotated with camera motion trajectories, object motion trajectories, and lighting directions, we construct three specialized datasets. All datasets consist of 25-frame video clips with a spatial resolution of $320\times 512$ pixels.
@@
Camera Motion Control Dataset. We construct this dataset from RealEstate10K [100], curating 62,000 clips with smooth camera trajectories. As shown in Fig. 3, for each clip, we take the first frame as the reference image, use DUSt3R to reconstruct a globally aligned 3D point cloud, and render the point cloud along the ground-truth camera trajectory to produce geometry-aware renderings. Since RealEstate10K lacks captions, we uniformly sample 4 frames from each clip and use Qwen2-VL-7B-Instruct [101] to generate a clip-level caption. Although RealEstate10K is predominantly indoors, the collected clips cover diverse camera motions, providing rich supervision for fine-grained camera control.
Figure 3: Construction pipeline of the Camera Motion Control Dataset. We reconstruct a 3D point cloud from the first frame with DUSt3R, render views along the ground-truth camera trajectory, and generate a clip-level caption with Qwen2-VL-7B-Instruc.
@@
Object Motion Control Dataset. We construct this dataset from WebVid-10M [1], consisting of 60,000 video clips. The dataset construction pipeline, illustrated in Fig. 4, follows these five steps: (1) Clip Filtering: Clips with abrupt scene changes are removed using PySceneDetect [^1] https://github.com/Breakthrough/PySceneDetect, and sequences with temporal intervals (1–16 frames) are sampled. To retain only clips with significant motion, optical flow is computed via MemFlow [102], and the bottom 25% of clips with low motion scores are filtered out. (2) Grid-point Tracking: For each filtered clip, we place a $16\times 16$ grid on the first frame and track each grid point across the clip using CoTrackerV3 [103]. (3) Dense Trajectory Sampling: For each trajectory, the average per-frame displacement is computed and normalized by the image diagonal. A clip is discarded as camera-dominated if at least 60% of trajectories exceed 3% of the diagonal; otherwise, it is retained as object-motion-only. Within retained clips, trajectories shorter than the clip-level average length are removed to ensure meaningful object motion. (4) Sparse Trajectory Sampling: From the filtered dense trajectories, 1–8 sparse trajectories are sampled with probability proportional to their length. (5) Optical Flow Smoothing: Optical flow between adjacent frames is computed to encode motion direction and intensity. Finally, a Gaussian filter is applied to smooth the sparse trajectory matrix, ensuring stable training.
Figure 4: Pipeline for the Object Motion Control Dataset: clip filtering, grid-point tracking, dense trajectory sampling, sparse trajectory sampling, and optical flow smoothing.
@@
Lighting Direction Control Dataset. Collecting videos that follow the same camera trajectory under different lighting conditions is impractical in real-world settings, making such data extremely difficult and costly to obtain. We introduce VideoLightingDirection (VLD), a synthetic dataset of 57,600 Blender-rendered videos across 3,600 static scenes, each providing 16 distinct lighting directions under an identical camera trajectory. As illustrated in Fig. 5, the construction pipeline comprises four steps: 1. Scene Creation. To better emulate real-world lighting scenarios, we design two complementary scene types: Haven and BOP. Haven uses Poly Haven HDR environment lighting together with a directional spotlight and places 3D models at the center of the HDR environment. BOP disables environment lighting, uses a single spotlight as the only light source, and randomly places BOP models in a six-plane textured room to simulate indirect lighting. Models and environments are randomly combined to increase diversity. (2) Camera Trajectory Sampling: We sample the starting position of the camera trajectory on a spherical shell centered at the 3D models with radius $r\in[0.7,1.3]$ meters. A smooth trajectory is then randomly generated around the model, ensuring that the camera always faces the center of the model. (3) Lighting Direction Sampling: To enhance the lighting effect, we reduce the HDR environment intensity by 40%. We uniformly sample 16 points on a hemisphere centered on the model, with the base surface normal aligned with the reference camera viewing direction. Each sampled point serves as the position of a 2kW spotlight (radius=1), oriented toward the model’s center. The lighting direction vectors are projected to the corresponding camera coordinate and normalized to obtain per-frame lighting direction labels (Sec. III-C). (4) Rendering and Annotation: Using Blender Cycles, we render each scene under 16 different lighting conditions while maintaining a consistent camera trajectory. Each frame is annotated with the camera trajectory and lighting direction, providing precise supervision signals for training.
Fig. 6 shows samples from the VideoLightingDirection (VLD) dataset. We include two types of scenes: (a) Haven scenes, in which 3D models from Poly Haven are placed centrally within HDR environments; and (b) BOP scenes, featuring randomly positioned BOP models in six-plane textured rooms to simulate indirect lighting. We display two lighting directions per scene, while the dataset actually provides 16 distinct lighting directions under an identical camera trajectory. This design highlights how lighting direction affects shading, reflections, and visual coherence, offering valuable data for training models with precise lighting annotations.
Figure 5: Construction pipeline for the VideoLightingDirection Dataset. A 3D model is randomly paired with an HDR environment to create a static scene; a smooth camera trajectory and a single spotlight are then sampled around the 3D model and rendered in Blender to produce a video clip with the lighting direction annotation.
III-E Training
We train VidCRAFT3 conditioned on text, a reference image, camera motion, object motion, and lighting direction, using a standard denoising diffusion objective. Since no single dataset provides all three control annotations jointly, we adopt a three-stage progressive training strategy.
Training Objective. Let $z_{0}$ denote the clean video latent of a training sample $x$, and let $c=\{c_{\text{img}},c_{\text{txt}},c_{\text{cam}},c_{\text{obj}},c_{\text{light}}\}$ denote the control signals. Here, $c_{\text{txt}},c_{\text{img}},c_{\text{cam}},c_{\text{obj}}$, and $c_{\text{light}}$ refer to text prompt, reference-image latent, point cloud renderings, multi-scale motion features, and lighting embedding, respectively. We train the noise estimator $\epsilon_{\theta}$ to reverse the diffusion process:
$\min_{\theta}\;\mathbb{E}_{z_{0},c,t,\epsilon\sim\mathcal{N}(0,I)}\!\left[\big\|\epsilon-\epsilon_{\theta}(z_{t},t,c)\big\|_{2}^{2}\right],$ (3)
where $\epsilon$ denotes Gaussian noise, $t\!\in\!\{0,\ldots,T\}$ is the diffusion timestep, and $z_{t}$ is the noised latent at step $t$. During training, we apply classifier-free dropout: with probability $p_{\text{uncond}}$, we discard one randomly chosen conditioning branch or discard all; if $c_{\text{cam}}$ is discarded, it is replaced by the reference-image latent repeated $F$ times to match the temporal length (following DynamiCrafter), whereas other dropped conditionings are replaced with zero tensors of matching shape.
Figure 6: Illustrations of samples from the proposed VideoLightingDirection (VLD) Dataset, featuring synthetic scenes designed to model complex light-object interactions. (a) Haven scenes, where 3D models from Poly Haven are placed at the center of HDR environments. (b) BOP scenes, with BOP models randomly positioned within six-plane textured rooms to simulate indirect lighting. Each subset includes video frames captured under two distinct lighting conditions for visualization, maintaining consistent camera trajectories.
@@ III-E Training
Stage 1: Camera Motion Control Training. The training begins by initializing the model with DynamiCrafter pre-trained weights. The model is then fine-tuned on camera motion control dataset for 40,000 iterations. We optimize only the UNet while freezing the other modules to align the VDM with camera motion. This stage establishes robust 3D scene understanding by integrating point cloud renderings and aligns the VDM with precise global camera motion, while maintaining temporal consistency.
@@
Stage 2: Dense Object Trajectories and Lighting Mixed Fine-tuning. We combine the Object Motion Control Dataset and VLD Datasets to create a comprehensive dataset annotated with camera motion, object motion, and lighting direction. This hybrid dataset enhances the ability of model to learn joint control over these three conditions. We use dense object trajectories obtained from the “Trajectory Filtering by Length” step of the Object Motion Control Dataset pipeline; these dense trajectories are then processed by the same Optical Flow Smoothing procedure used for sparse trajectories to yield stable dense motion fields. Dense object trajectories are incorporated to provide rich motion details, accelerating model convergence. To retain the camera control capabilities from stage 1, the temporal layers of the UNet remain frozen, while the spatial layers and newly introduced components, including ObjMotionNet, lighting cross-attention, and the MLP for lighting direction projection, are optimized for 20,000 iterations. This stage enables the model to simultaneously control all three conditions while ensuring global camera alignment.
@@
Stage 3: Sparse Object Trajectories and Lighting Mixed Fine-tuning. The same hybrid dataset from Stage 2 is reused, but dense trajectories are replaced with sparse trajectories to simulate real-world user interactions. Fine-tuning the same parameters as Stage 2, with 20,000 iterations. This stage forces the model to infer complex motion patterns from limited trajectory data, and the progressive shift from dense to sparse supervision enhances its ability to generalize to practical scenarios.
By progressively learning camera motion, object motion, and lighting direction controls across stages while strategically freezing layers to retain prior knowledge, the training strategy yields fine-grained, synergistic control over all three elements without requiring fully annotated multi-task data.
III-F Inference
During inference, given a reference image $I_{\text{ref}}$ and a text prompt, users may optionally input one control or multiple controls, including camera trajectory $E$, sparse object trajectories $\mathcal{T}$, and a lighting direction $L_{\text{ref}}$. We reconstruct a 3D point cloud from $I_{\text{ref}}$ using Image2Cloud and render it along the camera trajectory to obtain geometry-aware renderings $R$ for camera motion. The sparse object trajectories $\mathcal{T}$ are encoded by ObjMotionNet into multi-scale motion features. The reference lighting direction $\mathbf{L}_{\text{ref}}$ is SH encoded and projected via an MLP, then fused through the Spatial Triple-Attention Transformer. In addition, text and image embeddings modulate the U-Net via the same transformer. Consistent with training, any missing condition is replaced by its “null” signal: when camera control is absent, a stationary camera is equivalent to rendering the reference view $F$ times, and for efficiency we directly repeat the reference image latent across $F$ frames; when object or lighting control is absent, we input zero tensors with matching shapes. This enables each control to operate alone or in any combination. We sample with DDIM and classifier-free guidance to generate video.
IV Experiments
Figure 7: Qualitative comparison of camera motion control on RealEstate10K. The left column shows target camera trajectory visualizations; rows show ground-truth sequences (GT), CameraCtrl[7], MotionCtrl[6], CamI2V[33], and Ours. VidCRAFT3 follows the camera trajectory more faithfully, preserving layout and parallax with fewer artifacts.
Figure 8: Qualitative comparison of object motion control on WebVid-10M. The left column shows target object trajectories visualizations; rows show ground-truth sequences (GT), Image Conductor [55], Motion-I2V [58], and Ours. VidCRAFT3 maintains identity and background while adhering to specified object trajectories, producing fewer artifacts and jitter across challenging scenes.
IV-A Experimental Setup